Please show all work!!

A mechanic’s variable costs are given by the function V C = 5c 2 + 50c, where c is the number of cars repaired. Fixed costs are $20. If the mechanic repairs a total of 10 cars, what is the average variable cost? (a) $100 (b) $102 (c) $150 (d) $200 (e) $75
Answer #1

Answer

The correct answer is (a) $100.

Average Variable Cost(AVC) is the Total variable cost per unit of output.

Formula :

Average Variable Cost(AVC) = Variable Cost/ Quantity

Here quantity = c and Variable cost = VC = 5c2 + 50c

Hence, Average Variable Cost(AVC) = VC/c = (5c2 + 50c)/c = 5c + 50

If 10 cars are repaired i.e. c = 10, Average Variable cost = 5c + 50 = 5*10 + 50 = 100

Thus, If the mechanic repairs a total of 10 cars, the average variable cost is $100

Hence, the correct answer is (a) $100.